# The frame-dragging vector potential on galaxy scales from DM-only Newtonian $N$-body simulations

### v2

@@ Abstract
pubyear: 2025pagerange: The frame-dragging vector potential on galaxy scales from DM-only Newtonian $N$-body simulationsA.2
@@ 1 Introduction
This challenge becomes even more pressing when studying the evolution of galaxies, the Milky Way in particular. Reconstructing its history with microarcsecond accuracy – a feat now achieved by missions such as Gaia (prustiGaiaMission2016) – demands the application of similarly precise, fully GR models for highly accurate astrometry (see e.g. Klioner_2003; crostaRayTracingAnalytical2015d).
@@
Traditionally, investigations into the formation and dynamics of galaxies rely on hydrodynamical simulations, which are high-resolution zoom-in refinements of larger-scale cosmological $N$-body simulations. To accurately reconstruct their evolutionary history - marked by multiple mergers and interactions - one must consider the possible contributions of purely non-linear GR phenomena. These effects may in principle introduce non-trivial corrections that may influence the inferred dynamical properties of galaxies, hence their evolution. The leading order beyond-Newtonian effect that appears in GR in post-Newtonian type approximations is gravito-magnetism, a vector potential in the metric that is responsible for the dragging of inertial frames. Far from being limited to strong fields, i.e. rotating neutron stars and black holes, this effect is ubiquitous, and relevant also in weak fields.
@@
Frame-dragging caused by the Earth’s rotation, which is linked to the gravito-magnetic vector potential, has been successfully measured by everittGravityProbeB. More recently, the presence of a gravitational dragging has been suggested by crostaTestingCDMGeometrydriven2020 to drive the rotation of the spacetime, effectively reproducing the flat rotation curve of the Milky Way measured with Gaia DR2 (gaiacollaborationGaiaDataRelease2018). This result has been confirmed with Gaia DR3 (gaiacollaborationGaiaDataRelease2023) and poses serious challenges to standard models of galactic dynamics (beordoGeometrydrivenDarkmattersustainedMilky2024e; beordoExploringMilkyWay2024), since it may potentially signal the breakdown of Newtonian weak-field approximations when applied to extended objects like galaxies.
In any case, even if these effects turn out to be small, as is typically expected in perturbative approaches, they may still be observable and should no longer be neglected, given the increasing precision of current and forthcoming surveys. Furthermore, these quantities could help constrain models of dark matter, dark energy, and modified gravity theories, as different contributions are expected (thomasGravityNonlinearScales2015).
@@
Post-processing methods, such as those developed by bruniComputingGeneralrelativisticEffects2014a; thomasFullyNonlinearPostFriedmann2015, enabled for the first time the extraction of the gravito-magnetic vector potential from cosmological Newtonian $N$-body simulations on fully non-linear scales. This approach is justified within the post-Friedmann formalism (mililloMissingLinkNonlinear2015), where, at the leading order in the Newtonian dynamical regime, the purely relativistic vector potential is sourced by the transverse component of the momentum density field. While being about two or three orders of magnitude smaller than the scalar Newtonian potential, the vector potential computed from second-order perturbation theory (luCosmologicalBackgroundVector2009) is found to be underestimated down to two orders of magnitude at fully non-linear scales. Although the Newtonian simulations neglect feedback from the relativistic fields into the dynamics, these results confirm that at a sufficiently small scale, in the highly non-linear regime, the relativistic corrections to Newtonian dynamics are subdominant.
@@
With some simplifying approximations, GR N-body codes, such as gevolution (adamekNbodyMethodsRelativistic2014) and GRAMSES (barrera-hinojosaGRAMSESNewRoute2020) have emerged in recent years. Although these codes take into account the back reaction of the vector potential into the simulation dynamics, they lead to very similar findings (adamekGeneralRelativityCosmic2016; barrera-hinojosaVectorModesLCDM2021). Thanks to the Adaptive Mesh Refinement (AMR) technique, which enables higher spatial resolution to be achieved in high-density regions, barrera-hinojosaVectorModesLCDM2021 also quantified the gravito-magnetic acceleration inside dark matter halos, where the ratio with respect to the standard Newtonian acceleration remains typically of the order of $10^{-3}$ for redshift $z<3$ and halo mass in the range $10^{12.5}\text{--}10^{15}\;h^{-1}\,\mathrm{M}_{\odot}$.
@@
Based on a fluid-description but otherwise fully relativistic, codes such as the Einstein Toolkit (EinsteinToolkit) have also been used in cosmology in recent years (BentivegnaBruni; MunozBruni) to study possible GR effects. In particular, using idealised initial conditions based on a simple spatial distribution in the growing mode of a curvature perturbation, it was shown in MunozBruni that gravito-magnetic effects are always generated around halos and filaments. Current work based on the N-body relativistic code GRAMSES (barrera-hinojosaGRAMSESNewRoute2020) and assuming similar initial conditions confirms these effects (Barrerainprep).
@@
However, these studies focused on large cosmological simulations with box size $L>160\;h^{-1}$ Mpc and mass resolutions greater than $10^{8}\;h^{-1}\,\mathrm{M}_{\odot}$. While these configurations are well-suited for studying large-scale structure formation, they lack the necessary spatial and mass resolution to accurately capture the internal dynamics of galaxies. This work makes use of simulations from the IllustrisTNG suite (nelsonIllustrisTNGSimulationsPublic2019), which are well suited for resolving small-scale structures and are specifically designed for galactic studies. In future works, the use of such a suite will also straightforwardly facilitate the extension to the hydrodynamical counterpart of the simulations in order to investigate the role of baryons in this context, which has yet to be evaluated in the literature.
@@
To address this limitation, the post-Friedmann (PF) approach to cosmology has been developed by mililloMissingLinkNonlinear2015 as a new approximation scheme that provides a unified framework for all scales, from the fully non-linear Newtonian regime to the largest scales where relativistic effects become important (Bruni2012; Bruni_gaussianity2014). This approach generalises to a FLRW background spacetime the weak-field approximation on a Minkowski background, expanding Einstein field equations in inverse powers of the speed of light, following the pioneering work by PN_chandrasekhar (see PoissonWill, for a comprehensive modern treatment). However, while in the PN formalism velocities are treated as small, in the PF approach only peculiar velocities are small.
@@
When the set of equations is linearised, the formalism correctly reproduces first-order cosmological relativistic perturbation theory. On the other hand, at the leading order (which we refer to as 0PF, following mililloMissingLinkNonlinear2015) the Newtonian regime is recovered. Therefore, the scheme is capable of describing the evolution of structures on all scales of interest in a unified framework (mililloMissingLinkNonlinear2015). In the weak-field limit, however, besides the standard Newtonian scalar gravitational potential, a new vector gravitational potential appears as sourced by energy-momentum currents. This vector potential physically represents the gravito-magnetic field that manifests as the GR effect of frame-dragging.
@@ 2.1 The formalism
where $a$ and $\delta_{ij}$ are respectively the scale factor and flat spatial metric of the FLRW background [^1] Latin indices refer to the spatial coordinates and can take the values 11,22,33.. The metric perturbations are characterized by the scalar potentials $U$ and $V$, the vector potential $\mathbf{B}$ [^2] Note that the vector potential $\mathbf{B}$ defined in the metric tensor (1) of the PF formalism (with $B_{i}$ components) is defined with a different convention compared to barrera-hinojosaVectorModesLCDM2021, so that $B_{i}\equiv a\beta^{i}$., and the tensor potential $h_{ij}$. These perturbations are split into Newtonian (N) and post-Friedmann (P) quantities, each corresponding to a different level of approximation. Newtonian terms appear at orders $c^{-2}$ and $c^{-3}$ and are the only relevant ones at 0PF order, where EEs reduce to the exact non-linear equations of Newtonian cosmology for a pressure-less matter fluid. Post-Friedmann terms, appearing with powers $c^{-4}$ and $c^{-5}$ at 11PF order, contain the GR corrections modifying the Newtonian equations. The tensor perturbation $h_{ij}$ only enters the equations at 11PF order and represents pure tensor modes. The vector potential $\mathbf{B}$ and the tensor $h_{ij}$ only have two degrees of freedom each, since in the Poisson gauge the vector potential is divergence-less, namely $\nabla\cdot\mathbf{B}=0$, while $h_{ij}$ is transverse and trace-free.
@@
The Einstein field equations are then computed for the metric (1), including the cosmological constant. The equations for the conservation of energy and momentum are also expanded in powers of $1/c$ and calculated from the contracted Bianchi identities. We refer the reader to mililloMissingLinkNonlinear2015 for the full system of expanded equations obtained from the Einstein and hydrodynamic equations.
@@ 2.2 Leading order: frame-dragging vector potential in the Newtonian regime
where $\bar{\rho}$ denotes the background density and $\delta$ the density contrast, defined as $\delta\equiv(\rho-\bar{\rho})/\bar{\rho}$. Note that in this Newtonian regime, the density contrast is not required to be small.
Equation (6) coincides with the Poisson equation, where $V_{\mathrm{N}}$, usually identified as a curvature perturbation in relativistic perturbation theory (see e.g. Bruni_gaussianity2014), can be identified with the Newtonian gravitational potential generated by the matter field, $V_{\mathrm{N}}=-\phi_{\mathrm{N}}$. The scalar potential $U_{\mathrm{N}}$ enters in the Euler fluid equation (5), driving the dynamical flows of matter, and therefore can also be identified with the Newtonian gravitational potential, $U_{\mathrm{N}}=-\phi_{\mathrm{N}}$. The Einstein equations yield two additional constraints. Equation (7) forces the scalar potentials $V_{\mathrm{N}}$ and $U_{\mathrm{N}}$ to be equal, consistently with the fact that there is only one scalar potential in Newtonian theory, namely $V_{\mathrm{N}}=U_{\mathrm{N}}=-\phi_{\mathrm{N}}$. At higher orders in the $c^{-n}$ expansion, the difference between the two scalar potential is small but can be measured from relativistic $N$-body simulations (adamekGeneralRelativityCosmic2016).
The second constraint equation (8) relates the leading-order vector gravitational potential $B_{i}^{\mathrm{N}}$ to the momentum-density field of the matter. The latter is completely determined from the standard Newtonian equations (4) and (5), therefore the vector potential is not dynamical at the leading order, as it does not enter in the Euler equation, meaning that matter flow is not affected at this order of approximation. However, even in the regime where the matter dynamics is correctly described by Newtonian theory, the frame-dragging potential $B^{\mathrm{N}}_{i}$ should not be set to zero. If this were the case, an extra constraint on the Newtonian dynamics would be set. In fact, taking the the curl of Equation (8), the gravito-magnetic potential is sourced by the transverse part of the Newtonian momentum density $\mathbf{P}^{\mathrm{N}}$
@@
Even if the vector potential is not dynamical at this leading order, it appears in the metric and affects null geodesics and observables. For example, ThomasLensing found that the power spectra of the B-mode of shear for the weak gravitational lensing, which is generated by the time derivative of the vector potential, is $\approx 10^{-5}$ times the E-mode, which is the main effect generated by the scalar potential. barrera-hinojosaLookingTwistProbing2022 also showed that the possibility of detecting the gravito-magnetic effect in the lensing-kSZ cross-correlation is realistically possible with future CMB experiments.
The density and velocity fields sourcing the vector potential via Eq. (9) are purely Newtonian quantities and can be extracted from standard Newtonian $N$-body simulations (bruniComputingGeneralrelativisticEffects2014a; thomasFullyNonlinearPostFriedmann2015). Indeed, it is well known that transverse vector modes are generated by non-linearity, and in particular, vorticity is generated after shell crossing (pueblasGenerationVorticityVelocity2009; jelic-cizmekGenerationVorticityCosmological2018; hahnPropertiesCosmicVelocity2015). The extraction of the vector potential is illustrated in Section 4.

### v3

# The frame-dragging vector potential on galaxy scales from Dark-Matter-only Newtonian $N$-body simulations

@@ Abstract
pubyear: 2025pagerange: The frame-dragging vector potential on galaxy scales from Dark-Matter-only Newtonian $N$-body simulationsA.2
@@ 1 Introduction
This challenge becomes even more pressing when studying the evolution of galaxies, the Milky Way in particular. Reconstructing its history with microarcsecond accuracy – a feat now achieved by missions such as Gaia (Prusti et al., 2016) – demands the application of similarly precise, fully GR models for highly accurate astrometry (see e.g. Klioner_2003; Crosta et al., 2015).
@@
Traditionally, investigations into the formation and dynamics of galaxies rely on hydrodynamical simulations, which are high-resolution zoom-in refinements of larger-scale cosmological $N$-body simulations. To accurately reconstruct their evolutionary history – marked by multiple mergers and interactions – one must consider the possible contributions of purely non-linear GR phenomena. These effects may in principle introduce non-trivial corrections that may influence the inferred dynamical properties of galaxies, hence their evolution. The leading order beyond-Newtonian effect that appears in GR in post-Newtonian type approximations is gravito-magnetism, a vector potential in the metric that is responsible for the dragging of inertial frames. Far from being limited to strong fields, i.e. rotating neutron stars and black holes, this effect is ubiquitous, and relevant also in weak fields.
@@
Frame-dragging caused by the Earth’s rotation, which is linked to the gravito-magnetic vector potential, has been successfully measured by Everitt et al. (2011). More recently, the presence of a gravitational dragging has been suggested by Crosta et al. (2020) to drive the rotation of the spacetime, effectively reproducing the flat rotation curve of the Milky Way measured with Gaia DR2 (Gaia Collaboration et al., 2018). This result has been confirmed with Gaia DR3 (Gaia Collaboration et al., 2023) and poses serious challenges to standard models of galactic dynamics (Beordo et al., 2024a, b), since it may potentially signal the breakdown of Newtonian weak-field approximations when applied to extended objects like galaxies.
In any case, even if these effects turn out to be small, as is typically expected in perturbative approaches, they may still be observable and should no longer be neglected, given the increasing precision of current and forthcoming surveys. Moreover, the presence of a gravito-magnetic field affects not only the dynamics of matter but also photon trajectories, through which it may leave imprints on weak-lensing observables (schneiderGravitationalLenses1992). In this context, its contribution to the lensing convergence field has been found to be strongly suppressed relative to the dominant (standard) weak-lensing signal (Bartelmann:1999yn; Cuesta-Lazaro:2018uyv). Similarly, sagaWeakLensingInduced2015 showed that B-modes induced by this field remain below the expected noise level of galaxy surveys such as Vera Rubin LSST (LSST2019ApJ...873.. [^1] I). However, more recently, Barrera-Hinojosa et al. (2022) demonstrated that a cross-correlation method combining a weak-lensing survey with Cosmic Microwave Background (CMB) maps – particularly those containing the kinetic Sunyaev-Zel’dovich (kSZ) effect signal – can significantly enhance the signal-to-noise ratio compared to auto-correlation methods. This approach may therefore enable a successful detection with data from forthcoming CMB experiments such as the Simons Observatory (SimonsObservatory), in combination with galaxy surveys such as LSST and Euclid (euclid). Hence, as observational sensitivity improves, relativistic effects such as frame dragging may provide novel constraints on models of dark matter, dark energy, and modified gravity theories, since these scenarios predict distinct signals (Thomas et al., 2015a).
@@
Post-processing methods, such as those developed by Bruni et al. (2014); Thomas et al. (2015b), enabled for the first time the extraction of the gravito-magnetic vector potential from cosmological Newtonian $N$-body simulations on fully non-linear scales. This approach is justified within the post-Friedmann formalism (Milillo et al., 2015), where, at the leading order in the Newtonian dynamical regime, the purely relativistic vector potential is sourced by the transverse component of the momentum density field. Such vector potential, at the leading order, does not enter the Euler equation, so it does not affect matter dynamics. Therefore, even if the simulation is Newtonian, as it only accounts for the Newtonian scalar potential, the vector potential can be evaluated a posteriori to make a sensible estimate. While being about two or three orders of magnitude smaller than the scalar Newtonian potential, the vector potential computed from second-order perturbation theory (Lu et al., 2009) is found to be underestimated down to two orders of magnitude at fully non-linear scales. Although the Newtonian simulations neglect feedback from the relativistic fields into the dynamics, these results confirm that at a sufficiently small scale, in the highly non-linear regime, the relativistic corrections to Newtonian dynamics are subdominant.
@@
With some simplifying approximations, GR N-body codes, such as gevolution (Adamek et al., 2014) and GRAMSES (Barrera-Hinojosa and Li, 2020) have emerged in recent years. Although these codes take into account the back reaction of the vector potential into the simulation dynamics, they lead to very similar findings (Adamek et al., 2016; Barrera-Hinojosa et al., 2021). Thanks to the Adaptive Mesh Refinement (AMR) technique, which enables higher spatial resolution to be achieved in high-density regions, Barrera-Hinojosa et al. (2021) also quantified the gravito-magnetic acceleration inside dark matter halos, where the ratio with respect to the standard Newtonian acceleration remains typically of the order of $10^{-3}$ for redshift $z<3$ and halo mass in the range $10^{12.5}\text{--}10^{15}\;h^{-1}\,\mathrm{M}_{\odot}$.
@@
Based on a fluid-description but otherwise fully relativistic, codes such as the Einstein Toolkit (EinsteinToolkit) have also been used in cosmology in recent years (BentivegnaBruni; MunozBruni) to study possible GR effects. In particular, using idealised initial conditions based on a simple spatial distribution in the growing mode of a curvature perturbation, it was shown in MunozBruni that gravito-magnetic effects are always generated around halos and filaments. Current work based on the N-body relativistic code GRAMSES (Barrera-Hinojosa and Li, 2020) and assuming similar initial conditions confirms these effects (Barrerainprep).
@@
However, these studies focused on large cosmological simulations with box size $L>160\;h^{-1}$ Mpc and mass resolutions greater than $10^{8}\;h^{-1}\,\mathrm{M}_{\odot}$. While these configurations are well-suited for studying large-scale structure formation, they lack the necessary spatial and mass resolution to accurately capture the internal dynamics of galaxies. On the other hand, widely employed simulations suites like IllustrisTNG (Nelson et al., 2019) and EAGLE (The EAGLE team, 2017) are optimal for resolving small-scale structures and are specifically designed for galactic studies. The use of such simulations can also straightforwardly facilitate the extension to the hydrodynamical counterpart of the simulations in order to investigate the role of baryons in this context, which has yet to be evaluated in the literature.
The present work makes use of dark-matter-only simulations from the IllustrisTNG suite, following the post-processing method developed by Bruni et al. (2014) and Thomas et al. (2015b) to evaluate the gravito-magnetic potential in the post-Friedmann framework and its time evolution. This involves the use of the Delaunay Tessellation Field Estimator (DTFE, Schaap and van de Weygaert, 2000) to extract the continuous fields, and the spectral analysis for the solution of the Einstein’s equations. In the present paper, in addition, the solution over the entire simulation volume is also reconstructed from spectral analysis.
It is important to recall again that the vector potential is in all respects a relativistic quantity that has no counterpart in the conventional Newtonian dynamics, but only appearing at the leading order of the expansion in the post-Friedmann approach. Since it does not affect the dynamics at this order, the vector potential can be consistently estimated a posteriori from classical Newtonian N-body simulations, where the dynamics is only determined by the Newtonian scalar potential.
@@
To address this limitation, the post-Friedmann (PF) approach to cosmology has been developed by Milillo et al. (2015) as a new approximation scheme that provides a unified framework for all scales, from the fully non-linear Newtonian regime to the largest scales where relativistic effects become important (Bruni2012; Bruni_gaussianity2014). This approach generalises to a FLRW background spacetime the weak-field approximation on a Minkowski background, expanding Einstein field equations in inverse powers of the speed of light, following the pioneering work by PN_chandrasekhar (see PoissonWill, for a comprehensive modern treatment). However, while in the PN formalism velocities are treated as small, in the PF approach only peculiar velocities are small.
@@
When the set of equations is linearised, the formalism correctly reproduces first-order cosmological relativistic perturbation theory. On the other hand, at the leading order (which we refer to as 0PF, following Milillo et al., 2015) the Newtonian regime is recovered. Therefore, the scheme is capable of describing the evolution of structures on all scales of interest in a unified framework (Milillo et al., 2015). In the weak-field limit, however, besides the standard Newtonian scalar gravitational potential, a new vector gravitational potential appears as sourced by energy-momentum currents. This vector potential physically represents the gravito-magnetic field that manifests as the GR effect of frame-dragging.
@@ 2.1 The formalism
where $a$ and $\delta_{ij}$ are respectively the scale factor and flat spatial metric of the FLRW background [^1] Latin indices refer to the spatial coordinates and can take the values 11,22,33.. The metric perturbations are characterized by the scalar potentials $U$ and $V$, the vector potential $\mathbf{B}$ [^2] Note that the vector potential $\mathbf{B}$ defined in the metric tensor (1) of the PF formalism (with $B_{i}$ components) is defined with a different convention compared to Barrera-Hinojosa et al. (2021, with $\beta^{i}$ components), so that $B_{i}\equiv a\beta^{i}$., and the tensor potential $h_{ij}$. These perturbations are split into Newtonian (N) and post-Friedmann (P) quantities, each corresponding to a different level of approximation. Newtonian terms appear at orders $c^{-2}$ and $c^{-3}$ and are the only relevant ones at 0PF order, where the Einstein’s Equations (EEs) reduce to the exact non-linear equations of Newtonian cosmology for a pressure-less matter fluid. Post-Friedmann terms, appearing with powers $c^{-4}$ and $c^{-5}$ at 11PF order, contain the GR corrections modifying the Newtonian equations. The tensor perturbation $h_{ij}$ only enters the equations at 11PF order and represents pure tensor modes. The vector potential $\mathbf{B}$ and the tensor $h_{ij}$ only have two degrees of freedom each, since in the Poisson gauge the vector potential is divergence-less, namely $\nabla\cdot\mathbf{B}=0$, while $h_{ij}$ is transverse and trace-free.
@@
The Einstein field equations are then computed for the metric (1), including the cosmological constant. The equations for the conservation of energy and momentum are also expanded in powers of $1/c$ and calculated from the contracted Bianchi identities. We refer the reader to Milillo et al. (2015) for the full system of expanded equations obtained from the Einstein and hydrodynamic equations.
@@ 2.2 Leading order: frame-dragging vector potential in the Newtonian regime
where $\bar{\rho}$ denotes the background density and $\delta$ the density contrast, defined as $\delta\equiv(\rho-\bar{\rho})/\bar{\rho}$. Note that in this Newtonian regime, the density contrast is not required to be small. In the above equations, the lower script ",i,i" correspond to the partial derivative with respect to the spatial component $i$.
Equation (6) coincides with the Poisson equation, where $V_{\mathrm{N}}$, usually identified as a curvature perturbation in relativistic perturbation theory (see e.g. Bruni_gaussianity2014), can be identified with the Newtonian gravitational potential generated by the matter field, $V_{\mathrm{N}}=-\phi_{\mathrm{N}}$. The scalar potential $U_{\mathrm{N}}$ enters in the Euler fluid equation (5), driving the dynamical flows of matter, and therefore can also be identified with the Newtonian gravitational potential, $U_{\mathrm{N}}=-\phi_{\mathrm{N}}$. The Einstein equations yield two additional constraints. Equation (7) forces the scalar potentials $V_{\mathrm{N}}$ and $U_{\mathrm{N}}$ to be equal, consistently with the fact that there is only one scalar potential in Newtonian theory, namely $V_{\mathrm{N}}=U_{\mathrm{N}}=-\phi_{\mathrm{N}}$. At higher orders in the $c^{-n}$ expansion, the difference between the two scalar potential is small but can be measured from relativistic $N$-body simulations (Adamek et al., 2016).
The second constraint equation (8) relates the leading-order vector gravitational potential $B_{i}^{\mathrm{N}}$ to the momentum-density field of the matter. The latter is completely determined from the standard Newtonian equations (4) and (5), therefore the vector potential is not dynamical at the leading order, as it does not enter the Euler equation, meaning that matter flow is not affected at this order of approximation. However, even in the regime where the matter dynamics is correctly described by Newtonian theory, the frame-dragging potential $B^{\mathrm{N}}_{i}$ should not be set to zero. If this were the case, an extra constraint on the Newtonian dynamics would be set. In fact, taking the the curl of Equation (8), the gravito-magnetic potential is sourced by the transverse part of the Newtonian momentum density $\mathbf{P}^{\mathrm{N}}$
@@
Even if the vector potential is not dynamical at this leading order, it appears in the metric and affects null geodesics and observables. For example, ThomasLensing found that the power spectra of the B-mode of shear for the weak gravitational lensing, which is generated by the time derivative of the vector potential, is $\approx 10^{-5}$ times the E-mode, which is the main effect generated by the scalar potential. Barrera-Hinojosa et al. (2022) also showed that the possibility of detecting the gravito-magnetic effect in the lensing-kSZ cross-correlation is realistically possible with future CMB experiments.
The density and velocity fields sourcing the vector potential via Eq. (9) are purely Newtonian quantities and can be extracted from standard Newtonian $N$-body simulations (Bruni et al., 2014; Thomas et al., 2015b). Indeed, it is well known that transverse vector modes are generated by non-linearity, and in particular, vorticity is generated after shell crossing (Pueblas and Scoccimarro, 2009; Jelic-Cizmek et al., 2018; Hahn et al., 2015). The extraction of the vector potential is illustrated in Section 4.
At the next-to-leading order, 1PF, the correction $B^{\mathrm{P}}_{i}$ to the frame-dragging potential is of the order of $c^{-5}$ (see Equation 1b), meaning that its magnitude is $c^{2}$ times smaller than the leading order frame-dragging effect. In addition, at the 1PF expansion order, scalar and tensorial perturbations appear as $c^{-4}$ corrections to the metric terms and these, together with the frame-dragging, have a non-trivial effect on the matter dynamics, which is not investigated in this paper.